Circular No. 7359

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000E IN NGC 6951 R. Evans, Yerkes Observatory, University of Chicago; and G. Corso,

De Paul University, report that Cousins R and I CCD images taken on Jan. 24.03 UT with the 1-m Yerkes reflector show prediscovery images of SN 2000E (cf. IAUC 7351). Absolute photometry is not yet available, but preliminary analysis shows the supernova to be 2.7 mag brighter than the star USNO A2.0 1500.07508662 in R and 1.0 mag brighter in I.

Regarding their item on IAUC 7353, V. F. Polcaro et al. revise their information on the spectral features: “The spectra clearly show a number of narrow absorption features, overlapping the usual strong P-Cyg profiles typical of type-Ia supernovae. The stronger features are at 627.8 and 684.8 ± 0.3 nm, both with FWHM ∼ 0.5 nm.”

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD S. Lawrence and A. Crotts, Columbia University, report observations

confirming the new second hotspot (IAUC 7354) in the inner ring around SN 1987A: “Archival HST/STIS spectra taken on 1998 Nov. 14 show a spectral feature in Hα that is very similar to that in the first hotspot, arising from an unresolved source at radius 0′′.65 ± 0′′.03 and p.a. 103^{o} ± 5^{o} relative to the supernova. Note that this is 0′′.15 interior to the radius of maximum intensity radius of the ring. As might be expected from a ring/ejecta interaction at this location, the spot was redshifted by ∼ 100 km/s versus the ring material, with a velocity width of ∼ 100 km/s FWHM. The Hα flux was ∼ 1.5 × 10^{−15} erg s^{−1} cm^{−2}. No corresponding features are seen in forbidden lines.”

S. Maran, C. S. J. Pun, and G. Sonneborn, Goddard Space Flight Center, NASA, write: “Broadband optical images (FWHM = 0′′.1) obtained on Jan. 29 with the HST/STIS imaging mode (F28X50LP filter, 120-s exposure) confirm the recent brightening of a section of the inner circumstellar ring (cf. IAUC 7354). We find that the brightening is not confined to a single location, but is distributed from p.a. 90^{o} to 130^{o}. This result was determined by comparing the new images with images taken in 1998 November and 1999 February with the same instrumentation. We measure the brightest location to be at a distance of 0′′.65 in p.a. 126^{o} with respect to the supernova. The net count rate from this location, which appears as an unresolved source, is 3.1 ± 0.4 counts/s above the adjacent ring material. The brightest position is located near the inner edge of the ring, similar to the first ‘hotspot’ (cf. IAUC 6665, 6710, 6761, 7056).”

2000 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green